OpenStudy (anonymous):

What is the value of 3^2 over 3^4

OpenStudy (anonymous):

it has to be a fraction

OpenStudy (anonymous):

first simplify both exponents 3*3= ? 3*3*3*3=?

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

9 over 81?

OpenStudy (anonymous):

yes now simplify that fraction 9/81= ?

OpenStudy (anonymous):

oh 1/9 thanks!

OpenStudy (anonymous):

Please don't give direct answer @Martinaa its against the code of conduct and I was trying to help her understand what she was doing

OpenStudy (anonymous):

@Martinaa i know its 1/9.. i didnt know how to get there. thanks @kittenlover731

OpenStudy (anonymous):

your welcome just remember the steps 1. simplify the exponents 2. simplify the fraction

OpenStudy (anonymous):

alright thanks